COLLISION BROADENING OF NH_{3} BY D_{2} FROM 15 K TO 40 K

\underline{D. R. WILLEY}, R. E. TIMLIN, JR., I. SULAI, C. C. RUGGIERO, Dept. of Physics, Allegheny College, Meadville, PA 16335.

We report low temperature pressure broadening cross sections for the (J, K) = (1, 1), (2, 2) and (3, 3) inversion transitions of NH_{3} broadened by molecular deuterium from 15 to 40 K. These D_{2} broadening studies serve to complement our earlier work on the low temperature pressure broadening of NH_{3} by He, normal-H_{2} and para-H_{2}. While D_{2} has nearly the same mass as He, its electronic structure is similar to H_{2}. Owing to its greater mass and consequent decreased velocity relative to H_{2}, the expectation is that cross sections for NH_{3} collisions with D_{2} will be smaller than cross sections for H_{2} at a given temperature or energy. However, experimentally we find the NH_{3}-D_{2} cross sections are larger than for broadening by normal-H_{2}, J = 0, para-H_{2} and He.